# On the leptonic contribution to the ultra high-energy diffuse gamma-ray background

### v1

@@ 4 Discussion
In this Letter, we have investigated the contribution of unresolved sources to the Galactic diffuse gamma-ray flux. Our approach leads to a contribution in the inner Galaxy ranging from $\sim 20\%$ to $\sim 60\%$ of the flux of LHAASO around $10\,\rm{TeV}$, less than $\sim 40\%$ above $20\,\rm{TeV}$, and less than $\sim 20\%$ above $100\,\rm{TeV}$. It is always very small in the outer Galaxy.
Our results would still hold if a high-energy cutoff were to be introduced in eq. (2), as this would only decrease the already very small high-energy contribution of unresolved sources. In addition, we find that the approach described in section 2 leads to a spectrum for unresolved sources similar to that of resolved sources, confirming the self-consistency of our work.
Since our study is data-driven, we follow the actual, non-uniform age distribution of ATNF pulsars shown in figure 1. Another possibility considered by other authors, is to assume a uniform age distribution, with e.g. a birthrate of $0.7$ pulsars per century as in Dekker et al. (2024). However, doing so leads to model-dependent results substantially different from ours. Albeit not data-driven, we have also investigated this scenario by adding an artificial population of old pulsars to the data consistent with a constant Galactic pulsar birthrate of $\sim 5\,\rm{kyr}^{-1}$. We find that this would violate LHAASO observations. In particular, there would be $\sim 350-1300$ pulsars detectable by LHAASO in our simulations, which is clearly at odds with the results reported in the LHAASO catalog (Cao et al., 2024). In addition, we find that the vast majority of pulsars observed by Fermi at $\rm{GeV}$ energies in the vicinity of Earth (distance $\lesssim 4\,\rm{kpc}$) have a radio counterpart in ATNF and display the same age distribution as in ATNF for the same region, which further strengthens our choice to follow the ATNF distribution in age.
Our results are also in tension with those of Yan et al. (2024) except for the high energy part of the outer Galaxy. These tensions are due to the very different approaches adopted. We mostly rely on data avoiding as much as possible the inclusion of model-dependent theoretical hypotheses, while Yan et al. (2024) have adopted a more theoretical approach.
Finally, we note that our results are compatible with the Galactic neutrino flux of IceCube (Abbasi et al., 2023). Indeed, our $\sim 60\%$ upper limit on the leptonic contribution to the diffuse flux calls for a sizable hadronic contribution to it.
@@ 5 Conclusion
In conclusion, we have provided a new estimate of the contribution of unresolved leptonic sources to the Galactic diffuse gamma-ray background. Unlike previous studies which relied on theoretical assumptions, we adopted here a self-consistent data-driven approach so as to be model-independent. We notably provided stringent constraints on the contribution of unresolved leptonic sources at different energies in different parts of the sky. In particular, we find the leptonic contribution to be smaller than $60\%$ at $10\,\rm{TeV}$ and smaller than $20\%$ above $100\,\rm{TeV}$. This information allows to infer the true hadronic component of the diffuse background, whose characterization will be crucial to constrain cosmic-ray propagation, and ultimately unveil the nature of the long-sought-after sources of $\rm{PeV}$ cosmic rays.

### v2

@@ 4 Discussion
In this Letter, we have investigated the contribution of unresolved pulsars to the Galactic diffuse gamma-ray flux. Our approach leads to a contribution in the inner Galaxy ranging from $\sim 22\%\pm 9\%$ to $\sim 63\%\pm 17\%$ of the flux of LHAASO around $10\,\rm{TeV}$, less than $\sim 38\%\pm 10\%$ above $\sim 20\,\rm{TeV}$, and less than $\sim 21\%\pm 6\%$ above $100\,\rm{TeV}$. It is always below $\sim 18\%\pm 2\%$ in the outer Galaxy.
Our results would still hold if a high-energy cutoff were to be introduced in eq. (2), as this would only decrease the already very small high-energy contribution of unresolved pulsars. In addition, we find that the approach described in section 2 leads to a spectrum for unresolved pulsars similar to that of resolved pulsars, confirming the self-consistency of our work.
Since our study is data-driven, we follow the actual, non-uniform age distribution of ATNF pulsars shown in figure 1. Another possibility considered by other authors, is to assume a uniform age distribution, with e.g. a birthrate of $0.7$ pulsars per century as in Dekker et al. (2024). However, doing so leads to model-dependent results substantially different from ours. Albeit not data-driven, we have also investigated this scenario by adding an artificial population of old pulsars to the data consistent with a constant Galactic pulsar birthrate of $\sim 5\,\rm{kyr}^{-1}$. We find that this would violate LHAASO observations. In particular, there would be $\sim 350-1300$ pulsars detectable by LHAASO in our simulations, which is clearly at odds with the results reported in the LHAASO catalog (Cao et al., 2024). In addition, we find that the vast majority of pulsars observed by Fermi-LAT at $\rm{GeV}$ energies in the vicinity of Earth (distance $\lesssim 4\,\rm{kpc}$) have a radio counterpart in ATNF and display the same age distribution as in ATNF for the same region, which further strengthens our choice to follow the ATNF distribution in age.
Our results are also in tension with those of Yan et al. (2024) except for the high energy part of the outer Galaxy. In particular, the authors find that, in the inner Galaxy, unresolved pulsars account for more than 50% of the LHAASO flux in most of its energy range. Moreover, the fact that this contribution almost does not decrease with energy, suggests that its sepectrum is as hard as that of the diffuse background, which is in conflict with the steeper spectrum of the resolved LHAASO pulsars. These tensions are due to the very different approaches adopted. We mostly rely on data avoiding as much as possible the inclusion of model-dependent theoretical hypotheses, while Yan et al. (2024) have adopted a more theoretical approach, such as assuming a unique injection index $s=2.2$ for electron/positron pairs.
Finally, we note that our results are compatible with the Galactic neutrino flux of IceCube (Abbasi et al., 2023). Indeed, our $\sim 38\%\pm 10\%$ upper limit on the pulsar contribution to the diffuse flux above 202020\,20TeV calls for a sizable additional contribution to it, leaving ample room for a (large) hadronic component.
@@ 5 Conclusion
In conclusion, we have provided a new estimate of the contribution of unresolved pulsars, which are the most prominent UHE leptonic sources, to the Galactic diffuse gamma-ray background. Unlike previous studies which relied on theoretical assumptions, we adopted here a self-consistent data-driven approach. We notably provided stringent constraints on the contribution of unresolved pulsars at different energies in different parts of the sky. In particular, we find the pulsar contribution to be smaller than $\sim 38\%\pm 10\%$ at $20\,\rm{TeV}$ and smaller than $\sim 21\%\pm 6\%$ above $100\,\rm{TeV}$. While a contribution from other sources such as supernova remnants or micro-quasars cannot be ruled out, this information allows to infer robust upper limits on the hadronic component of the diffuse background, whose characterization will be crucial to constrain cosmic-ray propagation, and ultimately unveil the nature of the long-sought-after sources of $\rm{PeV}$ cosmic rays.